### v1

# Variants of a theorem of Macbeath in finite dimensional normed spaces

Abstract.
A classical theorem of Macbeath states that for any integers $d\geq 2$, $n\geq d+1$, $d$-dimensional Euclidean balls are hardest to approximate, in terms of volume difference, by inscribed convex polytopes with $n$ vertices. In this paper we investigate normed variants of this problem: we intend to find the extremal values of the Busemann volume, Holmes-Thompson volume, Gromov’s mass and Gromov’s mass of a largest volume convex polytope with $n$ vertices, inscribed in the unit ball of a $d$-dimensional normed space.
Key words and phrases:
theorem of Macbeath, polytope approximation, normed space, Busemann volume, Holmes-Thompson volume, Gromov’s mass, Gromov’s mass
2020 Mathematics Subject Classification:
52A21, 52A40, 52A27
Partially supported by the ERC Advanced Grant “ERMiD” and the National Research, Development and Innovation Office, NKFI, K-147544 grant.
1. Introduction
Approximation of a convex body by convex polytopes is a long-standing problem of convex geometry (see e.g. the surveys [5, 9, 14] or the book [13]). In this problem, given a convex body $K$, our goal is to find a member of a fixed family of polytopes with minimal ‘distance’ from the given body, i.e. which minimizes a given error function depending on the body and the polytope.
In this paper we investigate a classical variant of this problem: we intend to find convex polytopes contained in the body with minimal volume difference; that is, having maximal volume. One of the first results in this regard is possibly due to Blaschke [7, 49-53], who proved in 1923 that among unit area plane convex bodies, the area of a maximum area inscribed triangle is minimal for Euclidean disks. This result was generalized by Sas [26] in 1939, proving the same statement for inscribed convex $n$-gons for any fixed integer $n\geq 3$ in place of triangles. The $d$-dimensional case of this problem with $d\geq 3$ was settled by Macbeath [17], who, in a classical paper in 1957, proved that in any dimension, Euclidean balls are hardest to approximate, in terms of volume difference, by inscribed convex polytopes with $n$ vertices for any $n\geq d+1$. The goal of this paper is to find an analogue of this problem in finite dimensional normed spaces. We note that other variants of the theorem of Macbeath appeared in the recent manuscripts [6, 8], and that our approach to find normal variants of the problem of Macbeath was used also in [16] for a different problem.
[…]
Recall the well-known fact that any finite dimensional real normed space can be equipped with a Haar measure, and that it is unique up to multiplication of the standard Lebesgue measure by a scalar. Depending on the choice of this scalar, one may define more than one version of normed volume. We investigate the four variants of volume that are most regularly used in the literature. The Busemann and Holmes-Thompson volume of a set $S$ in a $d$-dimensional normed space with unit ball $B$, is defined as
(1.1) $\operatorname{vol}^{Bus}_{B}(S)=\frac{\kappa_{d}}{\lambda_{d}(B)}\lambda_{d}(S)\,\hbox{and}\,\operatorname{vol}^{HT}_{B}(S)=\frac{\lambda_{d}(B^{\circ})}{\kappa_{d}}\lambda_{d}(S),$
respectively. Note that the Busemann volume of the unit ball, and the Holmes-Thompson volume of its polar, are equal to that of a Euclidean unit ball. For Gromov’s mass, the scalar is chosen in such a way that the volume of a maximal volume cross-polytope, inscribed in the unit ball $B$ is equal to $\frac{2^{d}}{d!}$ , and for Gromov’s mass* (or Benson’s definition of volume), the volume of a smallest volume parallelotope, circumscribed about $B$, is equal to $2^{d}$. We denote the two latter quantities by $\operatorname{vol}^{m}_{B}(S)$ and $\operatorname{vol}^{m*}_{B}(S)$, respectively. For more information on these concepts of volume, the reader is referred to [3, 4].
Now we are ready to define the main problem investigated in this paper.
Definition 1.1.
For any $B\in\mathcal{K}^{d}_{o}$ and $\tau\in\{Bus,HT,m,m^{*}\}$, we set
(1.2) $\mu_{n}^{\tau}(B)=\{\operatorname{vol}_{B}^{\tau}(Q):Q\hbox{ is a polytope inscribed in }B\hbox{ with at most }n\textit{ vertices}\}.$
Problem 1.2.
For every $d\geq 2$, $n\geq d+1$ and $\tau\in\{Bus,HT,m,m^{*}\}$, find
$m^{\tau}(n,d)=\inf\left\{\mu_{n}^{\tau}(B):B\in\mathcal{K}^{d}_{o}\right\},\hbox{ and }M^{\tau}(n,d)=\sup\left\{\mu_{n}^{\tau}(B):B\in\mathcal{K}^{d}_{o}\right\}.$
Remark 1.3.
We observe that for any $d\geq 2$, $n\geq d+1$, $B\in\mathcal{K}^{d}_{o}$, nondegenerate linear transformation $L:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $n$, and $\tau\in\{Bus,HT,m,m^{*}\}$, $\mu_{n}^{\tau}(L(B))=\mu_{n}^{\tau}(B)$. Thus, by the compactness of affine classes of $o$-symmetric convex bodies (see [27]), the extremal values of the functionals in Problem 1.2 are attained at some convex bodies in $\mathcal{K}^{d}_{o}$.
In the paper, for any $B\in\mathcal{K}^{d}_{o}$, we denote by $I(B)$ and $C(B)$ a maximum volume cross-polytope inscribed in $B$, and a minimum volume parallelotope circumscribed about $B$, respectively. Furthermore, for any $n\geq d+1$, by $Q_{n}(B)$ we denote a maximum volume convex polytope in $B$ with at most $n$ vertices.
Remark 1.4.
For any $d\geq 2$, $n\geq d+1$ and $B\in\mathcal{K}^{d}_{o}$, we have
(1.3) $\begin{split}\mu_{n}^{Bus}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(B)}\kappa_{d},\quad\mu_{n}^{HT}(B)=\frac{\lambda_{d}(Q_{n}(B))\lambda_{d}(B^{\circ})}{\kappa_{d}},\\ \mu_{n}^{m}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(I(B))}\cdot\frac{2^{d}}{d!},\quad\mu_{n}^{m*}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(C(B))}\cdot 2^{d}.\end{split}$
The remaining part of the paper is structured as follows. In Section 2 we present our results about the Busemann volumes of inscribed polytopes. In Sections 3-5 we investigate the Holmes-Thompson volume, Gromov’s mass and Gromov’s mass of inscribed polytopes, respectively. Finally, in Section 6 we collect our additional remarks and open questions.
2. Extremal values of the Busemann volumes of inscribed polytopes
In this section we intend to find the values of $m^{Bus}(n,d)$ and $M^{Bus}(n,d)$ for $d\geq 2$ and $n\geq d+1$. In other words, we need to find the extremal values of
$\mu_{n}^{Bus}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(B)}\kappa_{d}.$
over all $B\in\mathcal{K}^{d}_{o}$. We start with some simple observations.
Remark 2.1.
By the result of Macbeath [17] mentioned in the introduction, for any $d\geq 2$ and $n\geq d+1$, $\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(B)}$ is minimal if $B=\mathbf{B}^{d}$. Thus, $m^{Bus}(n,d)=\mu_{n}^{Bus}(\mathbf{B}^{d})$ for all possible values of $d$ and $n$. From the result of Sas [26] it also follows that $m^{Bus}(n,2)=\mu_{n}^{Bus}(B)$ implies that $B$ is an ellipse.
Remark 2.2.
If $n\geq 2d$, then there are $o$-symmetric polytopes in $\mathbb{R}^{d}$ with at most $n$ vertices, implying that in this case $M^{Bus}(n,d)=\kappa_{d}$, and $M^{Bus}(n,d)=\mu_{n}^{Bus}(B)$ implies that $B$ is an $o$-symmetric polytope with at most $n$ vertices.
In the remaining part of this section we investigate $M^{Bus}(n,d)$ under the condition that $d+1\leq n\leq 2d-1$.
Remark 2.3.
Let $T$ be a triangle, and $o\in T$. Then an elementary computation shows that for $B=\operatorname{conv}(T\cup(-T))$, we have $\lambda_{2}(B)=2\lambda_{2}(T)$. On the other hand, $B$ is an $o$-symmetric convex hexagon if and only if $B=\operatorname{conv}(T\cup(-T))$ for some triangle $T$ satisfying $o\in T$. Thus, $M^{Bus}(3,2)=\frac{\pi}{2}$, and $\mu_{3}^{Bus}(B)=\frac{\pi}{2}$ if and only if $B$ is a (possibly degenerate) $o$-symmetric hexagon.
Our main result in this section is Theorem 2.4.
Theorem 2.4.
Let $d\geq 3$, with $d=4m+r$ for some integers $m,r$, where $0\leq r\leq 3$. Then
  • (a)
    $M^{Bus}(d+1,d)=\frac{\kappa_{d}}{\binom{d}{\lfloor\frac{d}{2}\rfloor}}$, and $M^{Bus}(d+1,d)=\mu_{n}^{Bus}(B)$ for $B\in\mathcal{K}^{d}_{o}$ if $B=\operatorname{conv}(S\cup(-S))$ for some regular simplex centered at $o$.
  • (b)
    If $0\leq r\leq 1$, then
    $M^{Bus}(d+2,d)=\frac{\kappa_{d}}{\binom{2m+1}{m}\cdot\binom{2m+r-1}{m}}.$
    Furthermore, $M^{Bus}(d+2,d)=\mu_{n}^{Bus}(B)$ for $B\in\mathcal{K}^{d}_{o}$ if $B=\operatorname{conv}(S_{1}\cup S_{2}\cup(-S_{1})\cup(-S_{2}))$, where $S_{1}$ and $S_{2}$ are $o$-symmetric regular simplices, of dimensions $12m+$ 1 and $12m+r-$ 1, respectively, lying in orthogonal linear subspaces;
  • (c)
    If $2\leq r\leq 3$, then
    $M^{Bus}(d+2,d)=\frac{\kappa_{d}}{\binom{2m+1}{m}\cdot\binom{2m+r-1}{m+1}}.$
    Furthermore, $M^{Bus}(d+2,d)=\mu_{n}^{Bus}(B)$ for $B\in\mathcal{K}^{d}_{o}$ if $B=\operatorname{conv}(S_{1}\cup S_{2}\cup(-S_{1})\cup(-S_{2}))$, where $S_{1}$ and $S_{2}$ are $o$-symmetric regular simplices, of dimensions $12m+$ 1 and $12m+r-$ 1, respectively, lying in orthogonal linear subspaces.
The proof of Theorem 2.4 is based on the properties of shadow systems (also called linear parameter systems), introduced by Rogers and Shephard [24] and studied by many authors.
Definition 2.5.
Let $X\subset\mathbb{R}^{d}$, let $v\in\mathbb{R}^{d}$ be a nonzero vector, and for any $x\in X$, let $\lambda_{x}\in\mathbb{R}$. Then the $1$-parameter family of sets
$C(t)=\operatorname{conv}\{x+t\lambda_{x}v:x\in X\},\quad t\in\mathbb{R}$
is called a shadow system.
We use the following result of Rogers and Shepard, stated as [24, Lemma 1].
Lemma 2.6.
With the notation in Definition 2.5, the function $t\mapsto\lambda_{d}(C(t))$ is convex.
To prove Theorem 2.4, we need some additional lemmas.
Lemma 2.7.
Let $\mathcal{H}$ be a family of finitely many hyperplanes in $\mathbb{R}^{d}$, where the intersection of all hyperplanes is $\{o\}$. Let $X=(x_{1},x_{2},\ldots,x_{k})\in\left(\mathbb{R}^{d}\right)^{k}$ be a finite sequence of points. Let $X_{H}$ denote the closure of the $k$-element point-sequences in $\left(\mathbb{R}^{d}\right)^{k}$ that can be obtained from $X$ by applying finitely many subsequent projections onto hyperplanes in $\mathcal{H}$ to each element of $X$ simultaneously. Then $(o,o,\ldots,o)\in X_{H}$.
Proof.
First, observe that since $X_{H}$ is closed and bounded, it is compact. In addition, for any projection $\operatorname{proj}_{H}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ onto a hyperplane $H\in\mathcal{H}$, we have $\operatorname{proj}_{H}(X_{H})\subseteq X_{H}$. Indeed, if $Y=(y_{1},y_{2},\ldots,y_{k})$ can be obtained from $X$ by finitely many orthogonal projections onto hyperplanes from $\mathcal{H}$, then the same is true about $\operatorname{proj}_{H}(Y)=(\operatorname{proj}_{H}(y_{1}),\operatorname{proj}_{H}(y_{2}),\ldots,\operatorname{proj}_{H}(y_{k}))$. Similarly, if $Y=(y_{1},y_{2},\ldots,y_{k})$ can be obtained as the limit of a sequence of such point sequences $Y^{m}=(y_{1}^{m},y_{2}^{m},\ldots,y_{k}^{m})$, then the same is true about $\operatorname{proj}_{H}(Y)$.
Let us define the function $f(Y)=\sum_{i=1}^{k}|y_{i}|$ on $\left(\mathbb{R}^{d}\right)^{k}$ for any sequence $Y=(y_{1},y_{2},\ldots,y_{k})$. This function is clearly continuous, and thus, it attains its minimum $\delta$ on $X_{H}$. If $\delta=0$, then we have $(o,o,\ldots,o)\in X_{H}$. Suppose for contradiction that $\delta>0$. Then, if $\sum_{i=1}^{k}|y_{i}|=\delta$ for the sequence $Y=(y_{1},y_{2},\ldots,y_{k})\in X_{H}$, we have that there is some $H\in\mathcal{H}$ that does not contain at least one of the $y_{i}$. But in this case the orthogonal projection $\operatorname{proj}_{H}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ onto $H$ strictly decreases the value of $f(Y)$, i.e. we have $f(\operatorname{proj}_{H}(Y))<f(Y)$, a contradiction. ∎
Lemma 2.8.
For any integers $d\geq 3$ and $1\leq k\leq\frac{d}{2}$, let
$A_{d}(k)=\binom{k}{\left[\frac{k}{2}\right]}\cdot\binom{d-k}{\left[\frac{d-k}{2}\right]},$
and let $A_{d}=\min\left\{A_{d}(k):1\leq k\leq\frac{d}{2}\right\}$. Let $d=4m+r$ for some $r\in\{0,1,2,3\}$. Then
  1. (1)
    for $r=0$, $A_{d}=\binom{2m-1}{m-1}\cdot\binom{2m+1}{m}$, with equality iff $k=2m-1$;
  2. (2)
    for $r=1$, $A_{d}=\binom{2m}{m}\cdot\binom{2m+1}{m}$, with equality iff $k=2m-1$ or $k=2m$;
  3. (3)
    for $r\in\{2,3\}$, $A_{d}=\binom{2m+1}{m}\cdot\binom{2m+r-1}{m+r-2}$, with equality iff $k=2m+1$.
Proof.
Consider the case that $k\leq\frac{d}{2}-2$. Then an elementary computation yields that
$A_{d}(k+2)=\frac{(k+2)\cdot(k+1)\cdot\lfloor\frac{d-k}{2}\rfloor\cdot\lceil\frac{d-k}{2}\rceil}{\lfloor\frac{k+2}{2}\rfloor\cdot\lceil\frac{k+2}{2}\rceil\cdot(d-k)\cdot(d-k-1)}\cdot A_{d}(k).$
Let $q_{d}(k)=\frac{A_{d}(k+2)}{A_{d}(k)}$. We show that $q_{d}(k)<1$ for all values of $k$ and $d$. We distinguish four cases depending on the parity of $k$ and $d$. If both $k,d$ are even, then
$q_{d}(k)=\frac{(k+1)(d-k)}{(k+2)(d-k-1)}.$
From this, $(k+1)+(d-k)=(k+2)+(d-k-1)$ and $0<k+1<k+2<d-k-1<d-k$ implies that $q_{d}(k)<1$. If $k$, $d$ or both are odd, a slight modification of this computation yields the same inequality.
Now, we prove the assertion. If $r=0$, then, by the previous consideration, $A_{d}(k)$ is minimal at $k=2m-1$ or $k=2m$. On the other hand,
$\frac{A_{d}(2m)}{A_{d}(2m-1)}=\frac{2m+2}{2m+1}>1,$
implying the statement in this case. In the cases $r=1$, $r=2$ and $r=3$ a similar computation proves the statement. ∎
Proof of Theorem 2.4.
The proof is based on a symmetrization procedure, similar to the one in [1], using shadow systems.
To prove (a), since the investigated quantity is affine invariant, we can clearly assume that $Q_{d+1}(B)$ is a translate of a regular simplex $S$ centered at the origin $o$, and $B=\operatorname{conv}\left((x+S\cup(-x-S)\right)$, where $x\in\mathbb{R}^{d}$.
Consider the function $f:\mathbb{R}^{d}\to\mathbb{R}$, $f(x)=\lambda_{d}(B)=\lambda_{d}(\operatorname{conv}\left((x+S\cup(-x-S)\right))$. Moving $x$ at a constant speed at any fixed direction induces a linear parameter system on the vertices of $B$. Thus, by Lemma 2.6, $f$ is a convex function. On the other hand, $f$ is invariant under any symmetry of $S$. In particular, it follows that the set of points where $f$ is minimal, is a convex set whose symmetry group contains that of $S$. Thus, $f$ is minimal at $o$. We note that the formula for the volume of $B$ in this special case can be found, e.g. in [25].
Now we deal with (b) and (c). Let $Q=Q_{d+2}(B)$. Then we may assume that $B=\operatorname{conv}(Q\cup(-Q))$. Using the characterization of $d$-dimensional convex polytopes with $d+2$ vertices (see e.g. [28] and the references therein), for some integer $11\leq k\leq d-$ 1, $Q$ can be written in the form $\operatorname{conv}\left((x_{1}+S_{1})\cup(x_{2}+S_{2})\right)$, where
  • (i)
    $S_{1}$ and $S_{2}$ are simplices centered at $o$, of dimensions $k$ and $d-k$, respectively;
  • (ii)
    the affine hulls of $x_{1}+S_{1}$ and $x_{2}+S_{2}$ intersect in a singleton, which is contained in both simplices.
Using the affine invariance of the problem, we may assume that $S_{1}$ and $S_{2}$ are regular simplices, and their affine hulls are orthogonal to each other. In this case the orthogonal projection of $Q$ onto the affine hull of $x_{1}+S_{1}$ is $x_{1}+S_{1}$, and the intersection of $Q$ with the affine hull of $x_{2}+S_{2}$ is $x_{2}+S_{2}$. Thus, by Lemma of [25], the volume of $Q$ is
$\lambda_{d}(Q)=\frac{1}{\binom{d}{k}}\lambda_{d}(S_{1})\lambda_{d}(S_{2}),$
independently of the position of the intersection point of $x_{1}+S_{1}$ and $x_{2}+S_{2}$ relative to $S_{1}$ and $S_{2}$. Let us define
$V(x_{1},x_{2})=\lambda_{d}\left(\operatorname{conv}\left((x_{1}+S_{1})\cup(x_{2}+S_{2})\cup(-x_{1}-S_{1})\cup(-x_{2}-S_{2})\right)\right).$
We intend to find the quantity
$V_{\min}(k)=\min\{V(x_{1},x_{2}):x_{1},x_{2}\in\mathbb{R}^{d},(x_{1}+S_{1})\cap(x_{2}+S_{2})\neq\emptyset\}.$
Note that $V_{\min}(k)\geq\min\{V(x_{1},x_{2}):x_{1},x_{2}\in\mathbb{R}^{d}\}$. Thus, to find $V_{\min}(k)$ it is sufficient to show that $V(x_{1},x_{2})$ is minimal over $x_{1},x_{2}\in\mathbb{R}^{d}$ if $x_{1}=x_{2}=o$. Let $H$ be the bisector of an edge of any of the simplices $S_{1}$ or $S_{2}$. Let $\operatorname{proj}_{H}:\mathbb{R}^{d}\to H$ denote the orthogonal projection of $\mathbb{R}^{d}$ onto $H$. Let $v$ denote a unit normal vector of $H$.
We show that $V(x_{1},x_{2})\geq V(\operatorname{proj}_{H}(x_{1}),\operatorname{proj}_{H}(x_{2}))$. To do it, we define a shadow system. Let the signed distance of $x_{1}$ and $x_{2}$ from $H$ be $\lambda_{1}$ and $\lambda_{2}$, respectively, where the sign is determined by $v$. For any $t\in\mathbb{R}$ and $i\in\{1,2\}$, set $x_{i}(t)=x_{i}+\lambda_{i}(t-1)v$, and
$K(t)=\operatorname{conv}\left((x_{1}(t)+S_{1})\cup(x_{2}(t)+S_{2})\cup(-x_{1}(t)-S_{1})\cup(-x_{2}(t)-S_{2})\right).$
Then $K(t)$ is a shadow system, and thus, the function $t\mapsto\lambda_{d}(K(t))$ is convex by Lemma 2.6. On the other hand, this function is an even function of $t$, as for any value of $t$, $K(-t)$ is the reflection of $K(t)$ to $H$. Thus, $\lambda_{d}(K(t))$ is minimal if $t=0$, implying, in particular, that $\lambda_{d}(K(1))\geq\lambda_{d}(K(0))$. But $\lambda_{d}(K(1))=V(x_{1},x_{2})$ and $\lambda_{d}(K(0))=V(\operatorname{proj}_{H}(x_{1}),\operatorname{proj}_{H}(x_{2}))$, which yields our statement.
Now, let $\mathcal{H}$ denote the family of the bisectors of all edges of $S_{1}$ and $S_{2}$. Note that the only common point of these hyperplanes is the origin $o$. Then, by Lemma 2.7, starting with the pair $(x_{1},x_{2})$, the pair $(o,o)$ belongs to the closure of the set of pairs that can be obtained by finitely many projections onto elements of $\mathcal{H}$. This, our previous argument, combined with the observation that $V(x_{1},x_{2})$ is a continuous function of $(x_{1},x_{2})$, yields that $V(x_{1},x_{2})\geq V(o,o)$ for any $x_{1},x_{2}\in\mathbb{R}^{d}$, or in other words, $V_{\min}(k)=V(o,o)$.
Next, we note that by [25], we have
$V(o,o)=\frac{1}{\binom{d}{k}}\lambda_{d}(\operatorname{conv}(S_{1}\cup(-S_{1})))\lambda_{d}(\operatorname{conv}(S_{2}\cup(-S_{2}))).$
Furthermore, again by [25], we have
$\lambda_{d}(\operatorname{conv}(S_{1}\cup(-S_{1})))=\binom{k}{\left[\frac{k}{2}\right]}\lambda_{d}(S_{1}),\hbox{ and}$
$\lambda_{d}(\operatorname{conv}(S_{2}\cup(-S_{2})))=\binom{d-k}{\left[\frac{d-k}{2}\right]}\lambda_{d}(S_{2}).$
Combining these inequalities, we obtain that for any $o$-symmetric convex body $K$ in $\mathbb{R}^{d}$, we have
$\mu_{d+2}^{Bus}(K)\leq\frac{\kappa_{d}}{\binom{k}{\left[\frac{k}{2}\right]}\cdot\binom{d-k}{\left[\frac{d-k}{2}\right]}}$
for some $11\leq k\leq d-$ 1. Thus the assertion in (2) and (3) of Theorem 2.4 follows from Lemma 2.8. ∎
3. Extremal values of the Holmes-Thompson volumes of inscribed polytopes
In this section, we intend to find the extremal values of
$\mu_{n}^{HT}(B)=\frac{\lambda_{d}(Q_{n}(B))\lambda_{d}(B^{\circ})}{\kappa_{d}}.$
over all $B\in\mathcal{K}^{d}_{o}$.
First, we collect our planar results.
Theorem 3.1.
We have the following.
  • (1)
    For any even integer $n\geq 4$, we have $M^{HT}(n,2)=\frac{n^{2}}{\pi}\sin^{2}\frac{\pi}{n}$, and $M^{HT}(n,2)=\mu_{n}^{HT}(B)$ for $B\in\mathcal{K}^{2}_{o}$ if $B$ is a regular $o$-symmetric $n$-gon.
  • (2)
    We have $m^{HT}(4,2)=\frac{6}{\pi}$, and $m^{HT}(4,2)=\mu_{4}^{HT}(B)$ for $B\in\mathcal{K}^{2}_{o}$ if $B$ is an affinely regular hexagon.
In the proof we need three lemmas. The first one, which is a generalization of a classical result of Dowker [11], appeared in [12].
Lemma 3.2.
Let $K\in\mathcal{K}^{2}_{o}$. Then, for every $m\geq 2$, among the maximum area $(2m)$-gons inscribed in $K$, there is an $o$-symmetric $(2m)$-gon.
For the next lemma, appearing as [22, Theorem 1] in a paper of Meyer and Reisner, recall that for every convex body $K\subset\mathbb{R}^{d}$, there is a unique point $s(K)\in\operatorname{int}(K)$, called the Santaló point of $K$, for which the quantity $\lambda_{d}\left((K-s(K))^{\circ}\right)$ is minimal. We note that if $K$ is $o$-symmetric, then we clearly have $s(K)=o$.
Lemma 3.3.
Let $C(t)$ be a shadow system as defined in Definition 2.5. Assume that for any $t\in[a,b]$, the set $C(t)$ has nonempty interior. Then the function
$t\mapsto\frac{1}{\lambda_{d}\left((C(t)-s(C(t)))^{\circ}\right)},t\in[a,b]$
is convex.
Finally, we need the following theorem of Coxeter [10] (for a more general version, see [15]).
Lemma 3.4.
Let $p_{1},\dots,p_{n}$ be the vertices of an $n$-gon $P$ in $\mathbb{R}^{2}$, in cyclic order. If there is some real number $\tau\geq 0$ such that $p_{j+2}-p_{j-1}=\tau(p_{j+1}-p_{j})$ for $j=1,\dots,n$, then $P$ is affinely regular.
In the proof we also need the concept of a Radon curve (see e.g. [20, 3, 16]). Such a curve is defined as the boundary of an $o$-symmetric convex disk with the property that for a linear image $B$ of this disk, $B^{\circ}$ coincides with the rotated copy of $B$ by $\frac{\pi}{2}$ around $o$. An equivalent characterization is given in [21]: the boundary of an $o$-symmetric convex disk $B$ is a Radon curve if and only if every point of $\operatorname{bd}(B)$ is the vertex of a largest area parallelogram inscribed in $B$.
Proof of Theorem 3.1.
First we prove (1). By Lemma 3.2, we may assume that $Q_{n}(B)$ is $o$-symmetric. Since $A\subseteq B$ implies $B^{\circ}\subseteq A^{\circ}$, we have $\mu_{n}^{HT}(B)\leq\mu_{n}^{HT}(Q_{n}(B))$. Thus, we may assume that $B=Q_{n}(B)$ is an $o$-symmetric $n$-gon. In this case $\mu_{n}^{HT}(B)$ is equal to the volume product of $B$, which is maximal if $B$ is an affinely regular convex $n$-gon (see [23], or [1] for an alternative proof).
Next, we prove (2). First, we find the minimum of $\mu_{4}^{HT}(B)$ in the family $\mathcal{P}_{o}^{k}$ of $o$-symmetric convex $k$-gons for an arbitrary fixed even integer $k\geq 6$. Note that by Lemma 3.2, we may assume that $Q=Q_{4}(B)$ is an $o$-symmetric parallelogram in $B$. First, note that by compactness, the minimum of $\mu_{4}^{HT}(B)$ is attained at some $o$-symmetric polygon $\hat{B}\in\mathcal{P}_{o}^{k}$ for any value of $k$. In addition, computing the value of $\mu_{4}^{HT}(B)$ for a parallelogram and a regular hexagon as $B$, we see that $\hat{B}$ has at least six sides. We show that to find the minimum of $\mu_{4}^{HT}(B)$ in $\mathcal{P}_{o}^{k}$ it is sufficient to consider the affinely regular $o$-symmetric $k^{\prime}$-gons for some even integer $6\leq k^{\prime}\leq k$.
Observe that if $\hat{B}$ has no vertex that is a vertex $p$ of some maximum area parallelogram inscribed in $\hat{B}$, then we may move $p$ and $-p$ in a direction that does not change $\lambda_{2}(Q_{4}(\hat{B}))$, but decreases $\lambda_{2}(\hat{B}^{\circ})$. Since in this case $\mu_{n}^{HT}(\hat{B})$ decreases, we conclude that every vertex of $\hat{B}$ is a vertex of at least one maximum area parallelogram inscribed in $\hat{B}$.
Let the vertices of $\hat{B}$ be $q_{1},q_{2},\ldots,q_{k^{\prime}}$, where $k^{\prime}\neq k$ is even. Consider a maximum area parallelogram $Q$ in $\hat{B}$, with vertices $q_{1},q_{i},-q_{1},-q_{i}$. Assume that $Q$ is the only maximum area parallelogram in $\hat{B}$ with $q_{1}$ as one of its vertices; this property is equivalent to the property that the supporting lines of $\hat{B}$ parallel to $[o,q_{1}]$ contain only $q_{i}$ and $-q_{i}$, respectively. Now we move $q_{1}$ and $-q_{1}$ symmetrically in a certain direction $v$. Note that by the linearity of directional derivatives, there is a unique direction $v$ such that if we move $q_{1}$ and $-q_{1}$ symmetrically in this direction, the derivative of $\lambda_{2}(\hat{B}^{\circ})$ is zero. It is easy to see that the line through $q_{1}$ parallel to $v$ is a supporting line of $\hat{B}$, intersecting it only at $q_{1}$. Thus, if $v$ is not parallel to $[o,q_{i}]$ then we can move $q_{1}$ and $-q_{1}$ to obtain an $o$-symmetric convex $k^{\prime}$-gon $B^{\prime}$ such that $\lambda_{2}(\hat{B}^{\circ})>\lambda_{2}({B^{\prime}}^{\circ})$ but $\lambda_{2}(Q_{4}(\hat{B}))=\lambda_{2}(Q_{4}(B^{\prime}))$. Since this is a contradiction, it follows, in particular, that the only maximum area parallelogram with $q_{i}$ as a vertex is $Q$.
Now, let us move $q_{1}$ as well as $-q_{1}$, to points $q_{1}^{\prime}$ and $-q_{1}^{\prime}$ parallel to $[o,q_{i}]$, until we reach some other sideline of $\hat{B}$, or $[o,q_{1}^{\prime}]$ is parallel to a side of $\hat{B}$ or $[q_{1}^{\prime},q_{2}]$ is parallel to $[o,q_{i+1}]$ or $[q_{k},q_{1}^{\prime}]$ is parallel to $[o,q_{i-1}]$. By Lemma 3.3, in this way we construct a an $o$-symmetric $k$-gon $B^{\prime}$ for which $\lambda_{2}(Q_{4}(\hat{B}))=\lambda_{2}(Q_{4}(B^{\prime}))$, and one of the following holds:
  • $\mu_{n}^{m}(\hat{B}^{\circ}))>\mu_{n}^{m}({B^{\prime}}^{\circ})$, or
  • $\mu_{n}^{m}(\hat{B}^{\circ})=\mu_{n}^{m}(B^{\prime\circ})$ and $B^{\prime}$ has strictly fewer vertices, or
  • $\mu_{n}^{m}(\hat{B}^{\circ})=\mu_{n}^{m}({B^{\prime}}^{\circ})$ and $B^{\prime}$ has strictly fewer vertices $q_{i}$ which belong to a unique maximum area parallelogram in $B^{\prime}$.
Since this process terminates after finitely many steps, we may assume that that for every vertex $q_{i}$ of $\hat{B}$, there is a side of $\hat{B}$ parallel to $[o,q_{i}]$. By the description of Radon curves before the proof of Theorem 3.1, it follows that $\operatorname{bd}(\hat{B})$ is a Radon curve. Thus, using a suitable linear transformation on $\hat{B}$, we may assume that $\hat{B}$ is the rotated copy of $\hat{B}$ by $\frac{\pi}{2}$ around $o$. By the construction method of Radon curves described in [21], we may assume that in this case the points (±$(\pm 1,0)$ and $(0,\pm 1)$ are the vertices of a largest area parallelogram in $\hat{B}$, implying also that $\lambda_{2}(Q_{4}(\hat{B}))=4$. Under these conditions, a simple counting argument shows that the number $k^{\prime}$ of the sides of $\hat{B}$ satisfies $k^{\prime}\equiv 2\mod 4$.
In the next part we deform $\hat{B}$ so that its boundary remains a Radon curve. We carry out symmetric deformations, that is, by modifying a vertex or a sideline of $\hat{B}$ in a certain way we always assume that the opposite vertex or sideline of $\hat{B}$ is modified accordingly.
Figure 1. An illustration for the modification of the Radon curve $\operatorname{bd}(\hat{B})$
Let us denote the side parallel to $[o,q_{1}]$ by $S_{i}=[q_{i},q_{i+1}]$, and let $x_{i}$ denote the intersection point of the sidelines through $S_{i-1}$ and $S_{i+1}$. Let us move the vertex $q_{1}$ continuously, parallel to $[q_{k^{\prime}},q_{2}]$ and at constant speed; we denote this moving vertex by $q_{1}(t)$, where $t$ stands for time and we set $q_{1}(0)=q_{1}$. By the properties of polarity, the points $q_{1}(t)$ correspond to lines through the point $w$ corresponding to the line of the points $q_{1}(t)$. Since $\hat{B}$ is the rotated copy of $\hat{B}$ by $\frac{\pi}{2}$ around $o$, $w$ is the rotated copy of the point $x_{i}$ by $-\frac{\pi}{2}$ around $o$, where $x_{i}$ is defined as the intersection point of $S_{i}$ and $[o,x_{i}]$. Based on this, let $S_{i}(t)$ be the closed segment the line through $x_{i}$, parallel to $q_{1}(t)$, with an endpoint $q_{i}(t)$ lies on the sideline through $S_{i-1}$, and the other endpoint $q_{i+1}(t)$ lying on the sideline through $S_{i+1}$ (see Figure 1). Now, let $\hat{B}(t)$ denote the $1$-parameter family of $o$-symmetric convex polygons obtained by replacing the vertices ±$\pm q_{1},\pm q_{i},\pm q_{i+1}$ of $\hat{B}$ by ±$\pm q_{1}(t),\pm q_{i}(t)$ and ±$\pm q_{i+1}(t)$, respectively. Then an elementary consideration shows that if $t$ is sufficiently small, then $\operatorname{bd}(\hat{B}(t))$ is a Radon curve. Clearly, for these values of $t$ we have $\lambda_{2}(Q_{4}(\hat{B}(t)))=4$. Note that $\lambda_{2}(\operatorname{conv}\{q_{k^{\prime}},q_{1},q_{2}\})=\lambda_{2}(\operatorname{conv}\{q_{k^{\prime}},q_{1}(t),q_{2}\})$. Thus, if $x_{i}$ is not the midpoint of $[q_{i},q_{i+1}]$, then for sufficiently small positive or negative values of $t$, we have $\lambda_{2}(\hat{B}(t))=\lambda_{2}(\hat{B}^{\circ}(t))<\lambda_{2}(\hat{B})=\lambda_{2}(\hat{B}^{\circ})$, contradicting our assumptions. Thus, we have that for every $i$, the midpoint of $[q_{i},q_{i+1}]$ lies on the segment $[o,x_{i}]$.
In the final part we show that the above properties imply that $\hat{B}$ is an affinely regular polygon. Using a suitable linear transformation, we can assume that $o,q_{i},q_{i+1}$ are vertices of an isosceles triangle. By the properties of Radon curves, this yields that the triangle $[q_{k},q_{1},q_{2}]$ is also isosceles, and the midpoint $z_{1}$ of $[q_{k},q_{2}]$ lies on $[o,q_{1}]$. Applying it for an arbitrary vertex we obtain that for any value of $j$, the midpoint $z_{j}$ of $[q_{j-1},q_{j+1}]$ belongs to $[o,q_{j}]$, and note that this property remains valid under any affine transformation.
Figure 2. An illustration for the proof that $\hat{B}$ is an affinely regular polygon.
For any $i$, let $L_{i}$ denotes the line through $[o,q_{i}]$, and set $\tau_{i}=\frac{|q_{i}|}{|z_{i}|}$. Using a suitable linear transformation and the fact that $z_{i}\in L_{i}$, we may assume that $q_{i-1},q_{i+1}$ are symmetric to $L_{i}$. Observe that $x_{i-1}$ can be obtained as the intersection of the line through $S_{i}$, and the line through $o$ and the midpoint of $S_{i-1}$. Similarly, $x_{i}$ can be obtained as the intersection of the line through $S_{i-1}$, and the line through $o$ and the midpoint of $S_{i}$. This yields that $x_{i-1}$ and $x_{i}$ are symmetric to $L_{i}$. Furthermore, since $z_{i-1}$ is the midpoint of $[q_{i-2},q_{i}]$, it follows that $q_{i-2}$ lies on the line $-q_{i}+L_{i-1}$, and thus, it is the intersection of $-q_{i}+L_{i-1}$ and the line through $[x_{i-1},q_{i-1}]$. Similarly, $q_{i+2}$ is the intersection of $-q_{i}+L_{i+1}$ and the line through $[x_{i},q_{i+1}]$. Hence, $q_{i-2}$ and $q_{i+1}$, as well as $z_{i-1}$ and $z_{i+1}$, are symmetric to $L_{i}$, which yields that $\tau_{i-1}=\tau_{i+1}$ (see Figure 2). On the other hand, since $\hat{B}$ is $o$-symmetric and the number of its vertices is not divisible by $4$, from this it follows that there is a value $\tau$ such that $\tau=\tau_{i}$ for all $i$. In other words, for every value of $i$, we have $2\tau q_{i}=q_{i+1}+q_{i-1}$. This yields that $2\lambda q_{i+1}=q_{i+2}+q_{i}$. Subtracting the first equality from the second one and rearranging the terms we obtain that
$(2\tau+1)(q_{i+1}-q_{i})=q_{i+2}-q_{i-1}$
for all values of $i$. Thus, by Lemma 3.4, $\hat{B}$ is an affinely regular polygon.
Finally, set $a_{m}=\mu_{4}^{HT}(P_{4m+2})$, where $m\geq 1$ and $P_{s}$ denotes a regular $s$-gon centered at $o$. An elementary computation shows that the sequence $\{a_{m}\}$ is increasing, and hence, it attains its minimum at $m=1$. ∎
We note that a straightforward modification of the proof of (a) of Theorem 2.4, replacing Lemma 2.6 in it with Lemma 3.3 readily yields the following.
Theorem 3.5.
Let $d\geq 3$. Let $B=\operatorname{conv}(S\cup(-S))$ for some simplex $S$ centered at the origin $o$. Then $M^{HT}(d+1,d)=\mu_{d+1}^{HT}(B)$.
4. Extremal values of the Gromov’s masses of inscribed polytopes
In this section, we intend to find the extremal values of
$\mu_{n}^{m}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(I(B))}\cdot\frac{2^{d}}{d!}.$
over all $B\in\mathcal{K}^{d}_{o}$.
First, we collect our planar results.
Theorem 4.1.
Let $n\geq 3$. Then the following holds.
  1. (1)
    We have $m^{m}(3,2)=1$, and $m^{m}(3,2)=\mu_{3}^{m}(B)$ for $B\in\mathcal{K}^{2}_{o}$ if $B$ is a parallelogram.
  2. (2)
    If $n\geq 4$, then $m^{m}(n,2)=2$, and $m^{m}(n,2)=\mu_{n}^{m}(B)$ for $B\in\mathcal{K}^{2}_{o}$ if $B$ is a parallelogram.
  3. (3)
    We have $M^{m}(4,2)=2$, and $M^{m}(4,2)=\mu_{n}^{m}(B)$ for any $B\in\mathcal{K}^{2}_{o}$.
  4. (3)
    If $n\geq 6$ is even, and $n^{\prime}$ denotes the largest positive integer with $n^{\prime}\leq n$ and $n^{\prime}\equiv 2\mod 4$, then $M^{m}(n,2)=\frac{n^{\prime}\sin\frac{2\pi}{n^{\prime}}}{2\sin\frac{8\pi}{n^{\prime}-2}}$, and $M^{m}(n,2)=\mu_{n}^{m}(B)$ if $B$ is an affinely regular $n^{\prime}$-gon.
Proof.
To prove (1), note that for every $B\in\mathcal{K}^{2}_{o}$,
$\mu_{3}^{m}(B)=\frac{\lambda_{d}(Q_{3}(B))}{\lambda_{d}(I(B))}\cdot 2\geq\frac{\lambda_{d}(Q_{3}(I(B)))}{\lambda_{d}(I(B))}\cdot 2=\frac{1}{2}\cdot 2=1,$
with equality e.g. if $I(B)=B$.
To prove (2), observe that if $n\geq 4$, then
$\mu_{n}^{m}(B)\geq\frac{\lambda_{2}(I(B))}{\lambda_{2}(I(B))}\cdot 2=2.$
The statement in (3) follows from Lemma 3.2. Now we prove (4).
Since $n$ is even, by Lemma 3.2, we may assume that $Q_{n}(B)$ is an $o$-symmetric convex $n$-gon. Thus,
$\mu_{n}^{m}(B)=\frac{2\lambda_{2}(Q_{n}(B))}{\lambda_{2}(I(B))}\leq\frac{2\lambda_{2}(Q_{n}(B))}{\lambda_{2}(I(Q_{n}(B)))},$
with equality if $B=Q_{n}(B)$. In other words, we may assume that $B$ is an $o$-symmetric convex $n$-gon. From now on, we follow the proof of (2) of Theorem 3.1. We only sketch the proof.
Let $\hat{B}\in\mathcal{K}^{2}_{o}$ be a maximizer of the functional $\mu_{n}^{m}(\cdot)$. If $\hat{B}$ has a vertex $q$ which is not a vertex of a maximum area parallelogram in $\hat{B}$, then moving $q$ and $-q$ we can increase $\lambda_{2}(Q_{n}(\hat{B}))=\lambda_{2}(\hat{B})$ while keeping $\lambda_{2}(I(\hat{B}))$ fixed, a contradiction. Thus, every vertex of $\hat{B}$ is a vertex of a largest area parallelogram in $\hat{B}$.
If $\hat{B}$ has a vertex which belongs to a unique maximal area parallelogram in $\hat{B}$, then we can construct an $o$-symmetric $n$-gon $B^{\prime}$ such that
  • $\mu_{n}^{m}(\hat{B})<\mu_{n}^{m}(B^{\prime})$, or
  • $\mu_{n}^{m}(\hat{B})=\mu_{n}^{m}(B^{\prime})$ and $B^{\prime}$ has strictly fewer vertices, or
  • $\mu_{n}^{m}(\hat{B})=\mu_{n}^{m}(B^{\prime})$ and $B^{\prime}$ has strictly fewer vertices which belong to a unique maximum area parallelogram.
Thus, we may assume that every vertex $q$ of $\hat{B}$ belongs to more than one maximum area parallelogram, implying that $\operatorname{bd}(\hat{B})$ is a Radon curve, and that the number $n^{\prime}$ of the sides of $\hat{B}$ satisfies $n^{\prime}\equiv 2\mod 4$.
Next, as in the proof of Theorem 3.1 and using the notation there, we show that if the vertices of $B_{n}$ are $q_{1},q_{2},\ldots,q_{n^{\prime}}$ in counterclockwise order, $S_{i}=[q_{i},q_{i+1}]$, and $x_{i}$ is the intersection point of the sidelines of $\hat{B}$ through $S_{i-1}$ and $S_{i+1}$, then the segment $[o,x_{i}]$ contains the midpoint of $S_{i}$. This, as we have seen in the proof of Theorem 3.1, implies that $\hat{B}$ is affinely regular.
Finally, an elementary computation shows that if $g_{k}=\mu_{4k+2}^{m}(P_{4k+2})$, where $P_{4k+2}$ is an $o$-symmetric regular $(4k+2)$-gon, then the sequence $\{g_{k}\}$ is increasing, implying (3) of Theorem 4.1. ∎
Remark 4.2.
We observe that a slight modification of the proof of (3) of Theorem 4.1 yields the following: If $n\geq 3$ is not necessarily even, then there is some $\hat{B}\in\mathcal{K}^{2}_{o}$ such that $\operatorname{bd}(\hat{B})$ is a Radon curve, and $M^{m}(n,2)=\mu_{n}^{m}(\hat{B})$.
Our higher dimensional results are collected in the next theorem. For this, we need the following straightforward consequence of Lemma 2.6, the proof of which we leave to the reader.
Lemma 4.3.
For any convex polytope $P$ in $\mathbb{R}^{d}$ and $n\geq d+1$, there is a maximum volume polytope $Q_{n}$ with at most $n$ vertices inscribed in $P$ such that every vertex of $Q_{n}$ is a vertex of $P$.
Theorem 4.4.
Let $d\geq 3$. We have the following.
  1. (1)
    For any $n\geq 2d$, $m^{m}(n,d)=\frac{2^{d}}{d!}$. Furthermore, $m^{m}(n,d)=\mu_{n}^{m}(B)$ for $B\in\mathcal{K}^{d}_{o}$ if $B$ is a cross-polytope.
  2. (2)
    For $d+1\leq n\leq 2d-1$, $m^{m}(n,d)=\frac{2^{d}}{d!\cdot 2^{2d-n}}$, and $m^{m}(n,d)=\mu_{n}^{m}(B)$ for $B\in\mathcal{K}^{d}_{o}$ if $B$ is a cross-polytope.
Proof.
To prove (1), observe that $\lambda_{d}(Q_{n}(B))\geq\lambda_{d}(I(B))$, with equality e.g. if $B=I(B)$. Thus,
$\mu_{n}^{m}(K)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(I(B))}\cdot\frac{2^{d}}{d!}\geq\frac{2^{d}}{d!},$
and the assertion in (1) follows.
To prove (2), note that
$\mu_{n}^{m}(B)\geq\frac{\lambda_{d}(Q_{n}(I(B)))}{\lambda_{d}(I(B))}\cdot\frac{2^{d}}{d!},$
and hence, we may assume that $B$ is a cross-polytope, that is $B=\operatorname{conv}\{\pm q_{i}:i=1,2,\ldots,d\}$ for some linearly independent vectors $q_{i}$. By Lemma 4.3, we may assume that every vertex of $Q_{n}(B)$ is a vertex of $B$. Since $Q_{n}(B)$ is $d$-dimensional, as it has positive volume, it contains at least one element of each pair $\{q_{i},-q_{i}\}$. Thus, the vertex set of $Q_{n}(B)$ consists of $(n-d)$ pairs of antipodal points, and $2d-n$ singletons. From this, an elementary computation yields the statement. ∎
5. Extremal values of the Gromov’s masss of inscribed polytopes
In this section, we intend to find the extremal values of
$\mu_{n}^{m*}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(C(B))}\cdot 2^{d}.$
over all $B\in\mathcal{K}^{d}_{o}$.
Our main result in this section is Theorem 5.1.
Theorem 5.1.
We have the following.
  1. (1)
    $M^{m*}(3,2)=2$, and $\mu_{3}^{m*}(B)=M^{m*}(3,2)$ for $B\in\mathcal{K}^{2}_{o}$ if $B$ is a parallelogram.
  2. (2)
    For any $n\geq 4$, $M^{m*}=4$, and $\mu_{n}^{m*}(B)=M^{m*}(n,2)$ for $B\in\mathcal{K}^{2}_{o}$ if $B$ is a parallelogram.
  3. (3)
    $m^{m*}(4,2)=2$, and $\mu_{4}^{m*}(B)=m^{m*}(4,2)$ for $B\in\mathcal{K}^{2}_{o}$ if $\operatorname{bd}(B)$ is a Radon curve.
  4. (4)
    If $n\geq 3$, then there is some $B\in\mathcal{K}^{2}_{o}$ such that $\mu_{n}^{m*}(B)=m^{m*}(n,2)$ and $\operatorname{bd}(B)$ is a Radon curve.
Proof.
To prove (1), note that
$\mu_{3}^{m*}(B)=\frac{4\lambda_{2}(Q_{3}(B))}{\lambda_{2}(C(B))}\leq\frac{4\lambda_{2}(Q_{3}(C(B)))}{\lambda_{2}(C(B))}=2.$
Similarly, to prove (2) observe that
$\mu_{n}^{m*}(B)\leq\frac{4\lambda_{2}(B)}{\lambda_{2}(B)}\leq 4,$
and we have equality if $B$ is a parallelogram.
For (3), we note that by Lemma 3.2:
$\mu_{4}^{m*}(B)=\frac{4\lambda_{2}(I(B))}{\lambda_{2}(C(B))}\geq 2,$
and we have equality if $\operatorname{bd}(B)$ is a Radon curve (see [21]).
Now we prove (4). Assume that $B\in\mathcal{K}^{2}_{o}$ satisfies $\mu_{n}^{m*}(B)=m^{m*}(n,2)$, and $B$ has minimum area among these convex disks. We show that $\operatorname{bd}(B)$ is an equiframed curve; i.e. every point of $\operatorname{bd}(B)$ lies on a side of a minimum area parallelogram circumscribed about $B$. Note that for any such parallelogram $P$, the midpoints of the sides of $P$ belong to $B$.
Suppose that there a supporting line $L$ of $B$ that is not a sideline of such a parallelogram, and note that by symmetry, $-L$ has the same property. Let $[p,q]=L\cap B$, where we possibly have $p=q$. If one of the endpoints, say $p$ is the midpoint of a side of a minimum area circumscribed parallelogram $P$, then replacing the sidelines of $P$ through $p$ and $-p$ by $L$ and $-L$ we again obtain a minimum area circumscribed parallelogram, which contradicts our assumption that $L$ is not a sideline of such a parallelogram. Thus, neither $p$ nor $q$ is the midpoint of a side of a minimum area circumscribed parallelogram. But then, slightly moving $L$ and $L^{\prime}$ towards $o$ we obtain some $o$-symmetric convex disk $B^{\prime}\subsetneq B$ with $\lambda_{2}(C(B^{\prime}))=\lambda_{2}(C(B))$, a contradiction. This shows that $\operatorname{bd}(B)$ is an equiframed curve.
Equiframed curves are investigated in [21], in particular, in Section 5 of [21] the authors describe a construction by which every equiframed curve can be constructed. Without detailing their construction, from that we just deduce the following statement: If $\operatorname{bd}(B)$ is an equiframed curve with $B\in\mathcal{K}^{2}_{o}$, then there is some $B^{\prime}\subseteq B$ in $\mathcal{K}^{2}_{o}$ such that $\operatorname{bd}(B^{\prime})$ is a Radon curve, and $\lambda_{2}(C(B^{\prime}))=\lambda_{2}(C(B))$, implying the statement. ∎
Remark 5.2.
Note that for any $n\geq d+1$ and $B\in\mathcal{K}^{d}_{o}$,
$\mu_{n}^{m*}(B)=\frac{\lambda_{d}(Q_{n}(B))}{\lambda_{d}(C(B))}\cdot 2^{d}\leq 2^{d}.$
On the other hand, if $n\geq 2^{d}$ and $B$ a parallelotope, then $Q_{n}(B)=B=P(B)$, implying that in this case $M^{m*}(n,d)=2^{d}$.
6. Additional remarks and open questions
The variant of Problem 1.2 for Busemann volume leads to the following question:
Problem 6.1.
Let $d\geq 2$ and $d+3\leq n\leq 2d-1$. Find the minimum volume of an $o$-symmetric convex polytope in $\mathbb{R}^{d}$ that contains a unit volume convex polytope with $n$ vertices.
For our next problem, note that since for every $B\in\mathcal{K}^{2}_{o}$, the sequence $\{\mu_{n}^{HT}(B)\}$ is increasing, the sequence $\{m^{HT}(n,2)\}$ is also increasing. On the other hand, the example of a parallelogram $P$ shows that for any $n\geq 4$,
$m^{HT}(n,2)\leq\frac{\lambda_{2}(P)\lambda_{2}(P^{\circ})}{\pi}=\frac{8}{\pi}.$
Observe that for any $B\in\mathcal{K}^{2}_{o}$, if $n$ is sufficiently large, then $\lambda_{2}(Q_{n}(B))\approx\lambda_{2}((B))$, implying that $\mu_{n}^{HT}(B)\approx\frac{\lambda_{2}(B)\lambda_{2}(B^{\circ})}{\pi}$. On the other hand, a classical result of Mahler shows that $\lambda_{2}(B)\lambda_{2}(B^{\circ})\geq 8$ for any $B\in\mathcal{K}^{2}_{o}$, with equality for parallelograms. This suggests that if $n$ is sufficiently large, then $m^{HT}(n,2)\approx\frac{8}{\pi}$.
We conjecture the following, which implies that $m^{HT}(n,2)=\frac{8}{\pi}$ for all $n\geq 7$.
Conjecture 1.
For any $B\in\mathcal{K}^{2}_{o}$, we have
$\lambda_{2}(Q_{6}(B))\lambda_{2}(B^{\circ})\geq 8,$
with equality if $B$ is a parallelogram.
We note that the ‘dual’ of this problem, stating that the quantity $\lambda_{2}(H(B))\lambda_{2}(B^{\circ})$, where $H$ is a smallest area hexagon circumscribed about $B$, is maximal for ellipses, is equivalent to a conjecture of Makai, Jr. [19] from 1978, which was proved very recently by Aliev [2].